c) The flow rate of a chocolate syrup is 4 m3s-1 and the cylindrical pipe in which the fluid flows has a diameter of 6 cm. Determine the Reynold’s number for the syrup which has a density of 1268 kgm-3 and a viscosity of 17 Pas. Indicate if the flow is laminar or turbulent
1
Expert's answer
2020-07-07T10:02:40-0400

The Reynold’s number for the cylindrical pipe is given by the following expression:


"Re = \dfrac{QD\rho}{A\mu}"

where "Q = 4m^3\/s" is the volumetric flow rate, "D = 0.06m" is the pipe diameter, "A = \pi D^2/4" is the pipe's cross-sectional area, "\rho = 1268kg/m^3" is a density and "\mu = 17Pa\cdot s" is is the dynamic viscosity.

Substituting this values to the equation, get:


"Re = \dfrac{4QD\rho}{\pi D^2\mu} = \dfrac{4Q\rho}{\pi D\mu} = \dfrac{4\cdot4\cdot 1268}{\pi \cdot 0.06\cdot 17} \approx 6331"

A flow can be considered as turbulent if "Re>2900". Thus, this flow is turbulent.


Answer. Re = 6331, turbulent.